# Algorithms and Complexity of Hedge Cluster Deletion Problems

### v1

@@ 6 Concluding remarks
In this paper, we introduced the Hedge Cluster Deletion problem. Our analysis for the complexity of the problem reveals that in almost every non-trivial underlying structure of a hedge graph the problem remains difficult. Aiming towards a poly(OPT)-approximation, we showed that the problem is Min Horn Deletion-complete. For this reason, we studied the hedge intersection graph. We have proposed a constant factor approximation algorithm whenever at most two hedges appear in any triangle. Moreover, for any acyclic hedge intersection graph we have given a polynomial-time algorithm. We believe that our results reveal an interesting path for future work on hedge optimization problems.
We conclude with some additional remarks and open problems related to our results. It is an open problem whether any Min Horn Deletion-complete problem is actually poly-APX-complete [KhannaSTW01, BliznetsCKP18]. The Min Horn Deletion-completeness of Hedge Cluster Deletion adds a natural edge deletion problem into the class of such unclassified problems. Moreover, interesting connections arise with respect to parameterized complexity. Recall that Hedge Cluster Deletion is easily shown to be FPT parameterized by the solution size. There seems to be a connection between the existence of polynomial kernel and poly(OPT)-approximation for a problem, because usually one can find poly(OPT)-approximation having polynomial kernel, and vice versa. Although this fact is not formal and there are problems that act differently [GiannopoulouLSS16], it is quite possible that such connection exists for edge deletion problems: up to now, the existence of poly(OPT)-approximation algorithms are similar to results for kernelization complexity [BliznetsCKP18, BelovaB22, CaiC15]. Therefore, it is natural to ask whether Hedge Cluster Deletion admits a polynomial kernel. Furthermore, an interesting point for future work is to consider structural parameters of the hedge intersection graph such as treewidth or cliquewidth. Our algorithm on acyclic hedge intersection graphs can be seen as a first step towards such an approach.

### v2

@@ 7 Concluding remarks
In this paper, we introduced the Hedge Cluster Deletion problem. Our analysis for the complexity of the problem reveals that in almost every non-trivial underlying structure of a hedge graph the problem remains difficult. Aiming towards a poly(OPT)-approximation, we showed that the problem is Min Horn Deletion-complete. From the parameterized perspective, we showed that the variant of the problem parameterized by the solution size does not admit polynomial kernel under reasonable hierarchy assumptions. On the positive side, we studied the hedge intersection graph. We have proposed a constant factor approximation algorithm whenever at most two hedges appear in any triangle. Moreover, for any acyclic hedge intersection graph we have given a polynomial-time algorithm. We believe that our results reveal an interesting path for future work on hedge optimization problems.
[…]
Theorem 7.1.
Propagational-$f$ Satisfiability is Min Horn Deletion-complete under $A$-reductions.
Proof.
By Theorem 4.7, it is enough to show $A$-reductions from Hedge Cluster Deletion to Propagational-$f$ Satisfiability and from Propagational-$f$ Satisfiability to Hedge Cluster Deletion. The latter is given in the ppt-reduction of Theorem 4.12 with the propagational function $f=f_{1}$, which serves as an $A$-reduction as well. Note that the described reduction is a polynomial-time reduction, hence an $A$-reduction, from Propagational-$f_{1}$ Satisfiability to Hedge Cluster Deletion, as the sizes of the corresponding solutions are exactly the same. For the other direction, we employ a proof similar to Lemma 4.6.
In particular, we show that there is an $A$-reduction from Hedge Cluster Deletion to Propagational-$f^{\prime}$ Satisfiability, where $f^{\prime}(x,y,z)=\lnot x\lor y\lor z$. It is not difficult to see that $f^{\prime}$ is indeed a propagational function (see also [CaiC15]). Let $H$ be a hedge graph given as an instance of Hedge Cluster Deletion, where every connected component of $H$ contains exactly three vertices by Lemma 2.5. We show how to construct a formula $\phi$ from $H$ such that every clause of $\phi$ is of the form $f^{\prime}(x,y,z)$. We correspond every hedge in $H$ to exactly one Boolean variable in $\phi$. For every $P_{3}$ or $K_{3}$ of $H$, we add a constant number of clauses in $\phi$ as explained below. Let $X,Y,Z$ be the hedges that appear in each component.
  • For every $K_{3}$ whose three edges belong to three distinct hedges $X,Y,Z$, we add three clauses $f^{\prime}(x,y,z),f^{\prime}(y,x,z),f^{\prime}(z,x,y)$ in $\phi$. It is not difficult to see that $f_{1}(x,y,z)=f^{\prime}(x,y,z)\land f^{\prime}(y,x,z)\land f^{\prime}(z,x,y)$, where $f_{1}$ is the function defined in Section 4.2.
  • For every $K_{3}$ where two of its edges belong to hedge $X$ and the other edge belongs to $Y$, we add the clause $f^{\prime}(y,x,0)$ in $\phi$. This can be translated as follows: whenever $y$ is set to $1$ it also enforces $x$ to $1$; otherwise, $y=0$ and $f^{\prime}(y,x,0)$ is satisfied, regardless of the assignment for $x$.
  • For every $P_{3}$ whose two edges belong to two distinct hedges $X,Y$, we add the clause $f^{\prime}(1,x,y)$ in $\phi$. This is justified by the fact that at least one of $x$ or $y$ must be set to $1$.
  • For every $P_{3}$ where both of its edges belong to the same hedge $X$, we add the clause $f^{\prime}(1,0,x)$ in $\phi$. As explained in 2.4, $x$ must be set to $1$.
Note that if there is a $K_{3}$ whose edges belong to a single hedge then we ignore such component. This completes the construction of $\phi$ which can be done in polynomial time. Similarly to the proof of Lemma 4.6, we conclude that for any subset $U$ of hedges in $H$, $H\setminus U$ is a cluster hedge-subgraph if and only if the one-assignment on the corresponding Boolean variables of $U$ satisfies $\phi$. Therefore, the polynomial-time transformation serves as an $A$-reduction, as claimed. ∎
Another interesting point for future work is to consider the sizes of the hedges (i.e., the number of edges in each hedge) instead of the total number of hedges, as a natural candidate for the objective of the problem. Then notice that within the corresponding extremities of the suggested objective the behavior is well-understood: if the number of edges in each hedge is large enough then the total number of hedges is rather small, which means that the problem can be solved efficiently; otherwise, if each hedge contains exactly one edge then the problem coincides with Cluster Deletion. This naturally suggests to exploit the complexity with respect to the variety on the sizes of the hedges. Furthermore, intriguing questions related to the complexity of Hedge Cluster Deletion arise when considering structural parameters of the hedge intersection graph such as treewidth or cliquewidth. Our algorithm on acyclic hedge intersection graphs can be seen as a first step towards such an approach.